Charge divide in two parts  and place at same finite distance of much 
 

Grade:12th pass

2 Answers

Arun
3 years ago
Dear Alok
 
Question is incomplete. Please fheck and rrpsotthis. You can also attach an image. I will be happy to help you.
Vikas TU
3 years ago
Deat student 
Let the charges be q and (Q – q). Suppose, the distance between them is ‘x’.
The force between them is, F = kq(Q – q)/x2 = (k/x2)(qQ – q2)
The force is extremum when dF/dq = 0
So,
d[(k/x2)(qQ – q2)]/dq = 0
=> (k/x2)(Q – 2q) = 0
=> q = Q/2
Thus, the force will be maximum when the charge on each object is Q/2.
Good Luck